# Analysis of a class of Recursive Distributional Equations Including the Resistance of the Series-parallel graph

### v1

@@ Abstract.
This paper analyzes a class of recursive distributional equations (RDE’s) proposed by Gurel-Gurevich [17] and involving a bias parameter $p$, which includes the logarithm of the resistance of the series-parallel graph. A discrete-time evolution equation resembling a nonlinear, fractional Fisher-KPP equation is derived to describe the CDF’s of solutions. When the bias parameter $p=\frac{1}{2}$, this equation is shown to have a PDE scaling limit, from which distributional limit theorems for the RDE are derived. Applied to the series-parallel graph, the results imply that $N^{-1/3}\log R^{(N)}$ has a nondegenerate limit when $p=\frac{1}{2}$, as conjectured by Addario-Berry, Cairns, Devroye, Kerriou, and Mitchell [1].
@@ 1.1. Resistance of the Series-Parallel Graph
The asymptotic analysis of $R^{(n)}$ as $n\to\infty$ was initiated in [18]. Since $G^{(n)}$ can be realized by combining two independent copies $G^{(n-1)}_{1}$ and $G^{(n-1)}_{2}$ of $G^{(n-1)}$, say, in series with probability $p$ and in parallel with probability 1p1-p, the law of the resistance $R^{(n)}$ is a solution of the recursive distributional equation (RDE)
@@
Using this fact, a number of results were proved in [18]: When $p>\frac{1}{2}$, the logarithm $\log R^{(n)}$ grows linearly in $n$. More precisely, there is a constant $\lambda(p)>0$ such that
(1) $\mathbb{P}\left\{(2p-1)\log 2\leq\frac{1}{N}\log R^{(N)}\leq\lambda(p)\right\}\to 1\quad\text{as}\,\,N\to\infty.$
Due to symmetry, this implies that $\log R^{(n)}\to$-$\infty$ linearly when $p<\frac{1}{2}$.
At the critical point $p=\frac{1}{2}$, Hambly and Jordan instead proved that $\log R^{(n)}$ exhibits sublinear growth. The main results of this paper (see Theorem 1 and Corollary 1 below) show that, in fact, $\log R^{(N)}$ grows like $N^{\frac{1}{3}}$. Specifially,
$\frac{1}{(72\zeta(3))^{\frac{1}{3}}N^{\frac{1}{3}}}\log R^{(N)}+\frac{1}{2}\overset{d}{\to}\text{Beta}(2,2)\quad\text{as}\,\,N\to\infty,$
@@
This paper endeavors to address this problem in some generality when $p=\frac{1}{2}$, focusing on the distributional limit of $n^{-1/3}X^{(n)}$. The general framework developed here applies not only to the resistance of the critical series-parallel graph, but also to some other RDE’s considered recently in the literature. For instance, as is shown in Section 4 below, the class of RDE’s determined by (2) and (5) includes the symmetric hipster random walk from [1] and two-player symmetric cooperative motion, one of the RDE’s analyzed very recently by Addario-Berry, Beckman, and Lin in [3].
While the analysis of the asymptotic behavior in the case $p\neq\frac{1}{2}$ is not touched on here, the framework developed in this paper suggests that it is related to the singular limit of certain fractional, quasilinear variants of the Fisher-KPP equation. This is made precise in Section 1.3.4 below.
@@ 1.2. General Setting
The joint law of the functions $(f_{+},f_{-})$ will be denoted by $\mathbf{P}$ in what follows. $\mathbf{P}$ is assumed to be a Borel probability measure on the space $BC([0,\infty))\times BC([0,\infty))$ of pairs of bounded continuous functions in $0,\infty)$ with the uniform norm topology. In most of the examples of interest here, however, $f_{+}$ and $f_{-}$ are deterministic; see the discussion in Section 4.
@@ 1.3. Evolution of the CDF
$F_{n+1}=TF_{n}.$
In the first step of this work, a discrete-time evolution equation is derived to describe the evolution of the CDF of $X^{(n)}$. Namely, there is a nonlinear operator $\mathscr{L}$ defined on the space of CDF’s such that the recursive equation $F_{n+1}=TF_{n}$ can be rewritten in the following suggestive form
(11) $F_{n+1}-F_{n}=\mathscr{L}F_{n}+(1-2p)F_{n}(1-F_{n}).$
@@
As far as convergence in distribution is concerned, deriving a limit for $n^{-1/\alpha}X^{(n)}$ for some exponent $\alpha$ is equivalent to proving a scaling limit for the equation (11). In the main result of this paper, which treats the critical point $p=\frac{1}{2}$, this is done by adapting the classical approach to numerical approximations of parabolic PDE’s developed by Barles and Souganidis [6].
@@
$F_{n+1}-F_{n}=|\nabla_{\mathbb{Z}}F_{n}|\Delta_{\mathbb{Z}}F_{n}\quad\text{in}\,\,\mathbb{Z}$,
where $\nabla_{\mathbb{Z}}$ and $\Delta_{\mathbb{Z}}$ are certain discrete derivative operators. This observation was then used in conjunction with the method of [6] and a regularization effect of the equation to show that, after a suitable rescaling, the CDF converges to the solution of the initial-value problem
@@
1.3.1. Assumptions
Some assumptions on the law $\mathbf{P}$ of the forcing $(f_{+},f_{-})$ are in order. First, define the class of functions $\mathcal{S}$ so that $f\in\mathcal{S}$ if and only if
@@ 1.3.1. Assumptions
Notice that the restriction of $g_{\rho}$ to the set $f(0),+\infty)$ is the right-continuous right-inverse of $u\mapsto u+f_{\rho}(u)$.
By definition, if $f_{\rho}(+\infty)=0$, then $g_{\rho}(s)-s\to 0$ as $s\to+\infty$. The next assumption asks that $g_{\rho}(s)-s$ vanishes in a quantitative manner:
(15) $\int_{0}^{\infty}(1+s)\mathbf{E}[|g_{+}(s)-s|+|g_{-}(s)-s|]\,ds<\infty.$
1.3.2. Main Result
Under the assumption (15), it is possible to define effective coefficients $\sigma\in\mathbb{R}$ and $a\geq 0$ according to the formulae
$\sigma$ $=-2p\int_{0}^{\infty}\mathbf{E}[|g_{+}(s)-s|]\,du+2(1-p)\int_{0}^{\infty}\mathbf{E}[|g_{-}(s)-s|]\,ds,\quad$
$a$ $=p\int_{0}^{\infty}\mathbf{E}[(3s-g_{+}(s))(s-g_{+}(s))]\,ds+(1-p)\int_{0}^{+\infty}\mathbf{E}[(3s-g_{-}(s))(s-g_{-}(s))]\,ds.$
The nonnegativity of $a$ follows from the fact that $g_{+}(s),g_{-}(s)\leq s$ for any $s\geq 0$.
Define the exponent $\alpha\in\mathbb{N}$ as follows:
$\alpha=2\quad\text{if}\,\,\sigma\neq 0,\quad\alpha=3,\quad\text{otherwise.}$
The next theorem characterizes the scaling limit of the CDF in the critical case $p=\frac{1}{2}$. The relevant background from PDE theory, particularly the theory of viscosity solutions, is reviewed in Appendix A below.
@@ 1.3.2. Main Result
Assume that $p=\frac{1}{2}$, $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$, and (15) holds. If $\{F_{n}\}$ is the sequence of CDF’s associated to some solution $\{X^{(n)}\}$ of (9), then, independent of the initial CDF $F_{0}$, if $F_{\epsilon}$ is the rescaled CDF defined by
$F_{\epsilon}(x,t)=F_{[\epsilon^{-\alpha}t]}(\epsilon^{-1}x),$
then $F_{\epsilon}\to F$ locally uniformly in $\mathbb{R}\times(0,\infty)$, where $F$ is determined as follows:
@@
    If $\sigma\neq 0$, then $F$ is the unique bounded discontinuous viscosity solution of the initial value problem
    (18) $\left\{\begin{array}[]{r l}\partial_{t}F-\sigma|\partial_{x}F|^{2}=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=\mathbf{1}_{[0,\infty)}(x).\end{array}\right$.
@@
    If $\sigma=0$, then $F$ is the unique bounded discontinuous viscosity solution of the initial-value problem
    (21) $\left\{\begin{array}[]{r l}\partial_{t}F-a|\partial_{x}F|\partial_{x}^{2}F=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=\mathbf{1}_{[0,\infty)}(x).\end{array}\right$.
These PDE’s become more familiar when passing to PDF’s: The spatial derivative $\rho=\partial_{x}F$ is a solution of Burger’s equation $\partial_{t}\rho=\sigma\partial_{x}(\rho^{2})$ in case (i) and a solution of the porous medium equation $\partial_{t}\rho=\frac{1}{2}a\partial_{x}^{2}(\rho^{2})$ in case (ii).
As observed in [3, 2], the solutions of the PDE’s above are explicitly known and closely related to the $\text{Beta}(2,1)$ and $\text{Beta}(2,2)$ distributions. The next corollary makes this precise.
Corollary 1.
Assume that $p=\frac{1}{2}$, $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$, and (15) holds, and let $\{X^{(n)}\}$ be a solution of the RDE (9) with an arbitrary initial distribution.
  • (i)
    If $\sigma\neq 0$, then
    $-\frac{\text{sgn}(\sigma)}{2\sqrt{|\sigma|}\sqrt{N}}X^{(N)}\overset{d}{\to}\text{Beta}(2,1).$
  • (ii)
    If $\sigma=0$ and $a>0$, then
@@
    If $\sigma=a=0$, then $\mathbf{P}\{f_{+}=f_{-}=0\}=1$ and $X^{(N)}\overset{d}{=}X^{(0)}$ for each $N$.
@@
In addition to characterizing the asymptotics of the resistance of the critical series-parallel graph, the proof of Theorem 1 also leads to a characterization of the limiting behavior of Pemantle’s min-plus binary tree (or first passage percolation on the critical series-parallel graph). This is discussed in Section 4.2 below.
1.3.3. Idea of Proof
@@
In the present context, monotonicity means the following: If $\{F_{n}\}$ and $\{G_{n}\}$ are two sequences of CDF’s arising from solutions of the RDE and if the inequality $F_{0}\leq G_{0}$ holds pointwise, then $F_{n}\leq G_{n}$ holds for any $n$. For the purposes of this paper, this is the bare minimum structural condition necessary to treat the evolution (11) as a parabolic equation.
Consistency, on the other hand, refers to the behavior of $T$ under rescaling. Specifically, Proposition 8 below shows that if $\Phi$ is a smooth CDF, rescaled according to $\Phi_{\delta}(x)=\Phi(\delta x)$, then
(22) $(\mathscr{L}\Phi_{\delta})(\delta^{-1}x)=\delta^{2}\sigma|\partial_{x}\Phi(x)|^{2}+\delta^{3}a|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)+\cdots.$
Naïvely inserting this expansion in the equation (11) and recalling the definition of the rescaled CDF $F_{\epsilon}$, one is led to guess that the PDE’s (18) and (21) ought to appear in the limit. The approach of [6] provides a framework for making such a formal argument rigorous.
1.3.4. Variants of the Fisher-KPP Equation when $p\neq\frac{1}{2}$
The expansion (22) leads to interesting new questions away from the critical point $p=\frac{1}{2}$. Passing to the ballistically rescaled CDF
$F_{\epsilon}(x,t)=F_{[\epsilon^{-1}t]}(\epsilon^{-1}x),$
one finds that formally
$\partial_{t}F_{\epsilon}$ $=\epsilon^{2}a|\partial_{x}F_{\epsilon}|\partial_{x}^{2}F_{\epsilon}+\epsilon\sigma|\partial_{x}F_{\epsilon}|^{2}+\frac{1}{\epsilon}(1-2p)F_{\epsilon}(1-F_{\epsilon})+\cdots.$
After discarding the error terms, this is the PDE obtained after applying a ballistic rescaling to the following quasilinear, fractional variant of the Fisher-KPP equation
(23) $\partial_{t}G=a|\partial_{x}G|\partial_{x}^{2}G+\sigma|\partial_{x}G|^{2}+(1-2p)G(1-G).$
The asymptotics of this equation, particularly the propagation of “front-like” solutions (such as CDF’s), does not seem to have been treated thus far in the literature. Both the cases where $a=0$ or $\sigma=0$ appear to be relevant to the problem at hand; see Section 4.3 for justification of this.

### v2

@@ Abstract.
This paper analyzes a class of recursive distributional equations (RDE’s) proposed by Gurel-Gurevich [20] and involving a bias parameter $p$, which includes the logarithm of the resistance of the series-parallel graph. A discrete-time evolution equation resembling a quasilinear Fisher-KPP equation is derived to describe the CDF’s of solutions. When the bias parameter $p=\frac{1}{2}$, this equation is shown to have a PDE scaling limit, from which distributional limit theorems for the RDE are derived. Applied to the series-parallel graph, the results imply that $N^{-1/3}\log R^{(N)}$ has a nondegenerate limit when $p=\frac{1}{2}$, as conjectured by Addario-Berry, Cairns, Devroye, Kerriou, and Mitchell [1].
@@ 1.1. Resistance of the Series-Parallel Graph
The asymptotic analysis of $R^{(n)}$ as $n\to+\infty$ was initiated in [21]. Since $G^{(n)}$ can be realized by combining two independent copies $G^{(n-1)}_{1}$ and $G^{(n-1)}_{2}$ of $G^{(n-1)}$, say, in series with probability $p$ and in parallel with probability 1p1-p, the law of the resistance $R^{(n)}$ is a solution of the recursive distributional equation (RDE)
@@
Using this fact, a number of results were proved in [21]: When $p>1/2$, the logarithm $\log R^{(n)}$ grows linearly in $n$. More precisely, there is a constant $\lambda(p)>0$ such that
(1) $\mathbb{P}\left\{(2p-1)\log 2\leq\frac{1}{N}\log R^{(N)}\leq\lambda(p)\right\}\to 1\quad\text{as}\,\,N\to+\infty.$
Due to symmetry, this implies that $\log R^{(n)}\to$-$\infty$ linearly when $p<1/2$.
At the critical point $p=1/2$, Hambly and Jordan instead proved that $\log R^{(n)}$ exhibits sublinear growth. The main results of this paper (see Theorem 2 and Corollary 1 below) show that, in fact, $\log R^{(N)}$ grows like $N^{1/3}$. Specifially,
$\frac{1}{(72\zeta(3))^{\frac{1}{3}}N^{\frac{1}{3}}}\log R^{(N)}+\frac{1}{2}\overset{d}{\to}\text{Beta}(2,2)\quad\text{as}\,\,N\to+\infty,$
@@
This paper endeavors to address this problem in some generality when $p=1/2$, focusing on the distributional limit of $n^{-1/3}X^{(n)}$. The general framework developed here applies not only to the resistance of the critical series-parallel graph, but also to some other RDE’s considered recently in the literature. For instance, as is shown in Section 4 below, the class of RDE’s determined by (2) and (5) includes the symmetric hipster random walk from [1] and two-player symmetric cooperative motion, one of the RDE’s analyzed very recently by Addario-Berry, Beckman, and Lin in [3].
While the analysis of the asymptotic behavior in the case $p\neq 1/2$ is not touched on here, the framework developed in this paper suggests that it may be amenable to techniques developed in the study of reaction-diffusion equations. This is discussed further in Section 1.5 below.
A few days after this paper first appeared as a preprint, a work of Chen, Duquesne, and Shi [13] appeared, which considers the same class of RDE’s and independently obtains the limit theorem for $n^{-1/3}X^{(n)}$.
@@ 1.2. General Setting
The joint law of the functions $(f_{+},f_{-})$ will be denoted by $\mathbf{P}$ in what follows. $\mathbf{P}$ is assumed to be a Borel probability measure on the space $BC([0,+\infty))\times BC([0,+\infty))$ of pairs of bounded continuous functions in $0,+\infty)$ with the uniform norm topology. In most of the examples of interest here, however, $f_{+}$ and $f_{-}$ are deterministic; see the discussion in Section 4.
@@ 1.3. Evolution of the CDF
$F_{n}=TF_{n-1}.$
In the first step of this work, this recursion is reformulated as a discrete-time evolution equation resembling a parabolic PDE. Namely, there is a nonlinear operator $\mathscr{L}$ defined on the space of CDF’s such that
(11) $F_{n}-F_{n-1}=\mathscr{L}F_{n-1}+(1-2p)F_{n-1}(1-F_{n-1}).$
@@
A key property of the equation (11) is its monotonicity: If $F\leq G$ pointwise, then $TF\leq TG$ also holds. This means that (11) is amenable to techniques developed in the theory of parabolic PDE’s. The monotonicity of $T$ also plays a fundamental role in the aforementioned work of Chen, Duquesne, and Shi [13]. Their proof involves controlling CDF’s through the delicate construction of explicit sub- and supersolutions of (11). By contrast, the proof here is more conceptual: It starts by establishing that the evolution equation (11) “looks” like a PDE under a certain rescaling of the variables $(x,n)$, then invokes abstract convergence results for monotone semigroups and parabolic PDE’s. In addition to the clear intuitive interpretation, the method has the advantage that it applies very generally, including to cases in which explicit formulae are not available.
1.4. Scaling Limits at the Critical Point $p=1/2$
As far as convergence in distribution is concerned, deriving a limit for $N^{-1/\alpha}X^{(N)}$ for some exponent $\alpha$ is equivalent to proving a scaling limit for the equation (11). In the main result of this paper, which treats the critical point $p=1/2$, this is done by adapting the classical approach to numerical approximations of parabolic PDE’s developed by Barles and Souganidis [7].
@@ 1.4. Scaling Limits at the Critical Point $p=1/2$
$F_{n}-F_{n-1}=|\nabla_{\mathbb{Z}}F_{n-1}|\Delta_{\mathbb{Z}}F_{n-1}\quad\text{in}\,\,\mathbb{Z}$,
where $\nabla_{\mathbb{Z}}$ and $\Delta_{\mathbb{Z}}$ are certain discrete derivative operators. This observation was then used in [3] in conjunction with the method of [7] and a regularization effect of the equation to show that, after a suitable rescaling, the CDF converges to the solution of the initial-value problem
@@
1.4.1. Main Results
The main result holds under some assumptions on the law $\mathbf{P}$ of $(f_{+},f_{-})$. First, define the class of functions $\mathcal{S}$ so that $f\in\mathcal{S}$ if and only if
@@ 1.4.1. Main Results
Notice that the restriction of $g_{\rho}$ to the set $f_{\rho}(0),+\infty)$ is the right-continuous right-inverse of the function $u\mapsto u+f_{\rho}(u)$.
By definition, if $f_{\pm}(+\infty)=0$, then $g_{\pm}(s)-s\to 0$ as $s\to+\infty$. The next assumptions ask that $g_{\pm}(s)-s$ vanishes in a quantitative manner: First, under the assumption that
(15) $\int_{0}^{+\infty}\mathbf{E}[|g_{+}(s)-s|+|g_{-}(s)-s|]\,ds<\infty,$
define the constant $\sigma$ by
$\sigma=-\int_{0}^{+\infty}\mathbf{E}[|g_{+}(s)-s|]\,du+\int_{0}^{+\infty}\mathbf{E}[|g_{-}(s)-s|]\,ds.$
If in addition the following stronger assumption holds
(16) $\int_{0}^{+\infty}(1+s)\mathbf{E}[|g_{+}(s)-s|+|g_{-}(s)-s|]\,ds,$
define also the constant $a\geq 0$ by
$a=\frac{1}{2}\int_{0}^{+\infty}\mathbf{E}[(3s-g_{+}(s))(s-g_{+}(s))]\,ds+\frac{1}{2}\int_{0}^{+\infty}\mathbf{E}[(3s-g_{-}(s))(s-g_{-}(s))]\,ds.$
The nonnegativity of $a$ follows from the fact that $g_{\pm}(s)\leq s$ for any $s\geq 0$.
The next two theorems characterize the scaling limit of the CDF at the critical point $p=1/2$ in terms of a family of parabolic PDE’s parametrized by $\sigma$ and $a$. The relevant background from PDE theory, particularly the theory of viscosity solutions, is reviewed in Appendix A below. Motivated by recent work of Chen, Derrida, Duquesne, and Shi [12], the theorem allows for the possibility that $p=p^{(N)}$ converges to $1/2$ as $N\to+\infty$.
@@
Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$, (15) holds, and $\sigma\neq 0$. Fix a $\theta\in\mathbb{R}$ and let $p^{(N)}=1/2+\theta N^{-1}$. For each $N\in\mathbb{N}$, let $\{X^{(N,n)}\}$ be a solution of the RDE
$X^{(N,n)}\overset{d}{=}\Phi^{(N)}(X^{(N,n-1)}_{1},X^{(N,n-1)}_{2})$,
where $(X^{(N,n-1)}_{1},X^{(N,n-1)}_{2})$ are i.i.d. copies of $X^{(N,n-1)}$ and $\Phi^{(N)}$ is given by (8) with bias parameter $p=p^{(N)}$.
Assume that there is a function $F_{\text{in}}$ such that, at the initial time $n=0$, the following limit holds:
$\lim_{N\to+\infty}\mathbb{P}\{N^{-1/2}X^{(N,0)}\leq x\}=F_{\text{in}}(x)\quad\text{for almost every}\,\,x\in\mathbb{R}.$
If $\{F_{N}\}$ is the sequence of rescaled CDF’s defined in $\mathbb{R}\times[0,+\infty)$ by
$F_{N}(x,t)=\mathbb{P}\{N^{-1/2}X^{(N,[Nt])}\leq x\},$
then $F_{N}\to F$ locally uniformly in $\mathbb{R}\times(0,+\infty)$ as $N\to+\infty$, where $F$ is the unique bounded discontinuous viscosity solution of the initial value problem
(19) $\left\{\begin{array}[]{r l}\partial_{t}F-\sigma|\partial_{x}F|^{2}+2\theta F(1-F)=0&\text{in}\,\,\mathbb{R}\times(0,+\infty),\\ F(x,0)=F_{\text{in}}(x).\end{array}\right$.
See Section 5.5 for an example in which (15) fails to hold and $N^{-1/2}X^{(N)}\overset{d}{\to}+\infty$.
If $\theta=\sigma=0$ above, then the proof of the theorem instead implies that $F_{N}$ converges to the initial datum $F_{\text{in}}$, meaning that $\{X^{(N,n)}\}$ exhibits no nontrivial motion in the diffusive scaling limit. Under the stronger assumption (16), one finds that instead the appropriate scaling is subdiffusive.
Theorem 2.
Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$, (16) holds, $\sigma=0$, and $a>0$. For each $N\in\mathbb{N}$, let $\{X^{(N,n)}\}$ be a solution of the RDE
$X^{(N,n)}\overset{d}{=}\Phi^{(N)}(X^{(N,n-1)}_{1},X^{(N,n-1)}_{2})$,
where $(X^{(N,n-1)}_{1},X^{(N,n-1)}_{2})$ are i.i.d. copies of $X^{(N,n-1)}$ and $\Phi^{(N)}$ is given by (8) with bias parameter $p=p^{(N)}$ satisfying $p^{(N)}=1/2+\theta N^{-1}$ for some $\theta\in\mathbb{R}$.
Assume that there is a function $F_{\text{in}}$ such that, at the initial time $n=0$, the following limit holds:
$\lim_{N\to+\infty}\mathbb{P}\{N^{-1/3}X^{(N,0)}\leq x\}=F_{\text{in}}(x)\quad\text{for almost every}\,\,x\in\mathbb{R}.$
If $\{F_{N}\}$ is the sequence of rescaled CDF’s defined in $\mathbb{R}\times[0,+\infty)$ by
$F_{N}(x,t)=\mathbb{P}\{N^{-1/3}X^{(N,[Nt])}\leq x\},$
then $F_{N}\to F$ locally uniformly in $\mathbb{R}\times(0,+\infty)$ as $N\to+\infty$, where $F$ is the unique bounded discontinuous viscosity solution of the initial-value problem
(22) $\left\{\begin{array}[]{r l}\partial_{t}F-a|\partial_{x}F|\partial_{x}^{2}F+2\theta F(1-F)=0&\text{in}\,\,\mathbb{R}\times(0,+\infty),\\ F(x,0)=F_{\text{in}}(x).\end{array}\right$.
The PDE’s derived above become more familiar when passing to PDF’s and setting $\theta=0$: The spatial derivative $\rho=\partial_{x}F$ is a solution of Burger’s equation $\partial_{t}\rho=\sigma\partial_{x}(\rho^{2})$ in Theorem 1 and a solution of the porous medium equation $\partial_{t}\rho=\frac{1}{2}a\partial_{x}^{2}(\rho^{2})$ in Theorem 2.
As observed in [3, 2], when $\theta=0$ and $F_{\text{in}}=\mathbf{1}_{[0,\infty)}$, the solutions of the PDE’s above are explicitly known and closely related to the $\text{Beta}(2,1)$ and $\text{Beta}(2,2)$ distributions. This leads to limit theorems for the RDE (9) started from an arbitrary (fixed) initial distribution.
Corollary 1.
Assume that $p=1/2$, $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$, and (15) holds, and let $\{X^{(n)}\}$ be a solution of the RDE (9) with an arbitrary initial distribution.
@@
    The sequence $\{N^{-1/2}X^{(N)}\}$ has the following distributional limit:
    $-\frac{\text{sgn}(\sigma)}{2\sqrt{N}}X^{(N)}\overset{d}{\to}\sqrt{|\sigma|}\text{Beta}(2,1),$
    where $\sqrt{|\sigma|}\text{Beta}(2,1)$ denotes a $\text{Beta}(2,1)$ random variable dilated by a factor $\sqrt{|\sigma|}$. In particular, if $\sigma=0$, then the limit is zero.
@@
    Under the stronger assumption (16), if $\sigma=0$ and $a>0$, then $\{N^{-1/3}X^{(N)}\}$ instead has a nontrivial distributional limit:
@@
    Under (16), if $\sigma=a=0$, then $\mathbf{P}\{f_{+}=f_{-}=0\}=1$ and $\{X^{(n)}\}$ is the constant sequence $X^{(n)}\overset{d}{=}X^{(0)}$.
@@
In addition to Corollary 1, Theorems 1 and 2 have two applications that are worth highlighting now. In Section 4.1 below, Theorem 2 is used to characterize the asymptotics of the resistance of the critical series-parallel graph in the general case when the edges have i.i.d. resistances, which could equal zero or $\infty$. Similarly, Section 4.2 gives a characterization of the limiting behavior of Pemantle’s min-plus binary tree (or first passage percolation on the critical series-parallel graph) when the weights on the leaves are i.i.d. with an arbitrary distribution, answering a question posed in [6].
1.4.2. Idea of the Proof
@@
As mentioned already above, monotonicity refers to the fact that the equation (11) is order-preserving: If $\{F_{n}\}$ and $\{G_{n}\}$ are two solutions, and if $F_{0}\leq G_{0}$ pointwise, then the inequality $F_{n}\leq G_{n}$ remains true for any subsequent $n$.
Consistency, on the other hand, refers to the behavior of the equation under rescaling. Specifically, Proposition 10 below shows that if $\Phi$ is a smooth CDF, rescaled according to $\Phi_{\delta}(x)=\Phi(\delta x)$, then
(23) $(\mathscr{L}\Phi_{\delta})(\delta^{-1}x)=\delta^{2}\sigma|\partial_{x}\Phi(x)|^{2}+\delta^{3}a|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)+\cdots\quad\text{as}\,\,\delta\downarrow 0.$
Rescaling the sequence of CDF’s $\{F_{n}\}$ according to $F_{N}(x,t)=F_{[Nt]}(N^{1/\alpha}x)$, naïvely inserting the above expansion into the equation (11), and setting $p=1/2+\theta N^{-1}$, one obtains the formal asymptotic expansion
$\partial_{t}F_{N}=N^{1-\frac{2}{\alpha}}\sigma|\partial_{x}F_{N}|^{2}+N^{1-\frac{3}{\alpha}}a|\partial_{x}F_{N}|^{2}-2\theta F_{N}(1-F_{N})+\cdots\quad\text{as}\,\,N\to+\infty,$
which motivates the choices $\alpha=2$ and $\alpha=3$ used above and the appearance of the PDE’s (19) and (22). The approach of [7] provides a framework for making such a formal argument rigorous.
1.5. Variants of the Fisher-KPP Equation when $p\neq\frac{1}{2}$
The results of this work suggest interesting new questions away from the critical point $p=\frac{1}{2}$. In view of the asymptotic expansion (23) and the PDE’s obtained in Theorems 1 and 2, the equation (11) shares some similarities with the reaction-diffusion equation
(24) $\partial_{t}G=a|\partial_{x}G|\partial_{x}^{2}G+\sigma|\partial_{x}G|^{2}-2\theta G(1-G).$
In fact, (11) becomes the discrete-in-space-time, finite-difference version of (24) for suitable choices of the forcing $(f_{+},f_{-})$; see Section 4.3 for more details.
A number of works in the PDE literature considered the equation (24) in the case when $\sigma=0$ and $a>0$. Using an auxiliary ODE, Enguiça, Gavioli, and Sanchez [17] established the existence of traveling waves propagating at any large-enough speed, as in the classical Fisher-KPP equation. The work of Audrito and Vázquez [5] establishes existence, uniqueness, and positivity properties of traveling waves and also analyzes the asymptotic behavior of solutions, proving that sufficiently rapidly decaying perturbations of the unstable state (say, 0 if $\theta<0$) converge to the stable state ($1$ if $\theta<0$) inside a ballistically growing region, with the rate of growth determined by the minimal wave speed. It would be interesting to extend such results to the present context.